# Transversal Structures in Graph Systems: A Survey

### v1

@@ 7.2 Transversal matching in hypergraph systems
Notice that the bound in Conjecture 7.8 is tight. In fact, one may consider the edge set of the complete $k$-graph on $(s+1)k-1$ vertices or the $k$-graph on $n$ vertices in which every edge intersects a fixed set of $s$ vertices. Setting $s=1$ in Conjecture 7.8, then it corresponds to the well-known Erdős–Ko–Rado theorem [45]. There are numerous papers devoted to the Erdős matching conjecture. For recent development on this conjecture, see [10, 54, 64, 89].
Many scholars have attempted to generalize the Erdős matching conjecture to transversal version in hypergraph systems. Matsumoto and Tokushige [112] generalized the Erdős–Ko–Rado theorem to two families, that is,
@@
It is natural to extend Erdős–Ko–Rado theorem to more than two hypergraphs. For this, Aharoni and Howard [6] and Huang, Loh and Sudakov [70] proposed the following conjecture independently.
@@
Let =$\mathcal{H}=\{H_{1},\ldots,H_{t}\}$ be a system of $k$-graphs on a common vertex set $V$ of size $n$. If $|E(H_{i})|>\max\left\{{\binom{n}{k}}-{\binom{n-t+1}{k}},{\binom{kt-1}{k}}\right\}$ for all $1\leq i\leq t$, then $\mathcal{H}$ has a transversal $t$-matching.
When $k=2$, this conjecture was confirmed by Meshulam (see [6]). Indeed, the case $k=2$ of this conjecture follows directly from an earlier result established by Akiyama and Frankl [8], as restated in [53]. Huang, Loh, and Sudakov [70] proved that Conjecture 7.10 holds for $n>3k^{2}t$. Keller and Lifshitz [84] proved that Conjecture 7.10 holds when $n\geq f(t)k$ for some large constant $f(t)$ which only depends on $t$, and this was further improved the bound on $n$ to $12kt\log(e^{2}t)$ by Frankl and Kupavskii [56]. Lu, Wang and Yu [108] improved this bound to $n>2kt$ for sufficiently large $t$. By using the sharp threshold techniques developed in [79], Keevash, Lifshitz, Long and Minzer [78] proved a more general result with $n=\Omega(kt)$, which is the optimal order of magnitude. Kupavskii [98] gave the concrete dependencies on the parameters by showing the conjecture holds for $n>3ekt$ with $t>10^{7}$. Recently, Gao, Lu, Ma and Yu [57] showed that Conjecture 7.10 is true when $k=3$ and $n$ is sufficiently large.
@@
Let =$\mathcal{H}=\{H_{1},\ldots,H_{t}\}$ be a $k$-partite $k$-graph systems with $n$ vertices in each partition class. If $|E(H_{i})|\geq(t-1)n^{k-1}$ for all $i\in[t]$, then $\mathcal{H}$ admits a transversal $t$-matching.
Aharoni and Howard proved that Conjecture 7.11 holds when $t=2$ [6] or $k\leq 3$ [7]. Lu and Yu [109] proved Conjecture 7.11 for $n>3(k-1)(t-1)$. Kiselev and Kupavskii [88] showed that this conjecture is true for $t\geq t_{0}$ where $t_{0}$ is a constant.
The degree conditions for the existence of large matchings in uniform hypergraphs have been extensively studied. Kühn, Osthus and Treglown [97] (also see Khan [85]) and Khan [86] determined the minimum $1$-degree condition for an $n$-vertex $3$-graph and 4-graph containing a perfect matching, respectively. Lu, Yu and Yuan [110] and Lu, Wang and Yu [107] proved the transversal version of the above results.
@@
Let $n$ be a sufficiently large integer and $k\in\{3,4\}$ with $k|n$. If =$\mathcal{H}=\{H_{1},\ldots,H_{\frac{n}{k}}\}$ is a $k$-graph system of order $n$ with $\delta_{1}(\mathcal{H})>\binom{n-1}{k}-\binom{(k-1)n/k}{k}$, then $\mathcal{H}$ contains a transversal perfect matching.
Recently, Lu and Wang [106] determined the minimum $1$-degree condition that guarantees a $3$-partite $3$-graph system admits transversal perfect matchings. This generalized a result of Lo and Markström [105] on the $1$-degree threshold for the existence of perfect matchings in $3$-partite $3$-graphs. Moreover, in their proof, they used fractional transversal matching theory obtained by Aharoni, Holzman and Jiang [5].
@@
Aharoni, Holzman and Jiang [5] studied the existence of rainbow fractional matching in a $k$-graph system.
@@
Let $k\geq 2$ be an integer, and let $n$ be a positive rational number. Assume that =$\mathcal{H}=\{H_{1},\ldots,H_{\lceil kn\rceil}\}$ is a $k$-graph system with $\nu^{*}(H_{i})\geq n$ for all $i\in[\lceil kn\rceil]$. Then $\mathcal{H}$ has a rainbow fractional matching of size $n.$
When the $k$-graph is $k$-partite and $n$ is an integer, the number of graphs in $\mathcal{H}$ required in Theorem 7.13 goes down from $kn$ to $kn-k+1.$ Theorem 7.13 provided a fractional version of the corresponding problem concerning rainbow matchings, which was solved by Drisko [42] and by Aharoni and Berger [1] for bipartite graphs (the rainbow matching problem remains open for general graphs as well as for $k$-partite hypergraphs when $k>2$).
Alon, Frankl, Huang, Rödl, Ruciński and Sudakov [9] established the relationship between perfect matching and perfect fractional matching by showing that every $n$-vertex $k$-graph $H$ with $\delta_{d}(H)\geq(\max\{c_{k,d},1/2\}+o(1))\binom{n-d}{k-d}$ has a perfect matching, and this condition is asymptotically best possible. Cheng, Han, Wang, and Wang [27] proved that the $d$-degree condition that ensures the existence a perfect matching in a $k$-graph also guarantees the existence of a transversal perfect matching in a $k$-graph system for $d\in[k-1]$.
@@
For every $\varepsilon>0$ and integer $d\in[k-1]$, there exists $n_{0}\in\mathbb{N}$, such that the following holds for all integers $n\geq n_{0}$ and $n\in k\mathbb{N}$. Every $n$-vertex $k$-graph system =$\mathcal{H}=\{H_{1},\ldots,H_{\frac{n}{k}}\}$ with $\delta_{d}(\mathcal{H})\geq(\max\{c_{k,d},\frac{1}{2}\}+\varepsilon)\binom{n-d}{k-d}$ contains a transversal perfect matching.
@@ 8 Further results
  • The rainbow Turán problem can be viewed as a special case of Turán problems in graph systems. Let $H$ be a graph. The rainbow Turán number of $H$, denoted by $\mathrm{ex}^{*}(n,H)$, is defined to be the the maximum number of edges in a properly edge-colored $n$-vertex graph $G$ with no rainbow copies of $H$. Notice that this properly edge-colored graph $G$ can be viewed as system $\mathcal{G}=\{G_{1},\ldots,G_{\chi(G)}\}$ of graphs where each $G_{i}$ is the graph consisting of edges with color $i.$ Then find a rainbow copy of $H$ in $G$ is equivalent to find a transversal copy of $H$ inside $\mathcal{G}$. The rainbow Turán number of some classical families of graphs were studied by Keevash, Mubayi, Sudakov and Verstraëte [80]. See more results in [17, 36, 73, 87].
  • In section 3, we only survey results on the minimum degree condition for transversal spanning structures in graph systems. In fact, some studies have considered the problem of embedding transversal non-spanning structures in a graph system. Li, Li and Li [101] studied the existence of transversal cliques in graph systems under Dirac-type conditions. Falgas-Ravry, Markström and Räty [49] studied the minimum degree version of Problem 2.2, and they found that the extremal behavior in the minimum degree condition differs strikingly from that seen in the density condition, with discrete jumps as opposed to continuous transitions.
  • The adjacency matrix of $G$ is $A(G)=(a_{ij})$, where $a_{xy}=1$ if $xy\in E(G)$ and 00 otherwise. The largest eigenvalue of $A(G)$ is called the spectral radius of $G$. The maximum spectral radius of graphs in a graph system that forbids transversal perfect matching, Hamilton path or Hamilton cycle was studied in [61, 67, 134].
  • One may consider the problem of embedding a rainbow graph $H$ in a graph system $\mathcal{G}$ provided that $H$ and all graphs in $\mathcal{G}$ belong to the same class. For example, Aharoni, Briggs, Holzman and Jiang [2] proved that every family of $12\left\lceil n/2\right\rceil-$ 1 odd cycles of order $n$ contains a rainbow odd cycle. For more results, we refer the reader to [41, 60, 100].
@@ 9 Conclusion
Although we have made an effort to provide a systematic coverage of recent classical results in graph systems, it is inevitable that some topics were left out of this survey due to constraints of space and time. Nevertheless, we would like to believe that we have presented enough results demonstrating how one can generalize known results in extremal graph theory to transversal version. Therefore we hope this survey will inspire further research in this area.
10 Acknowledgments
The authors would like to thank Yangyang Cheng and Jie Hu for helpful comments and valuable suggestions.

### v2

@@ 7.2 Matching and factor
Notice that the bound in Conjecture 7.8 is tight. In fact, one may consider the edge set of the complete $k$-graph on (s+1)k1(s+1)k-1 vertices or the $k$-graph on $n$ vertices in which every edge intersects a fixed set of $s$ vertices. Setting $s=1$ in Conjecture 7.8, then it corresponds to the well-known Erdős–Ko–Rado theorem [47]. There are numerous papers devoted to the Erdős matching conjecture. For recent development on this conjecture, see [10, 56, 67, 95].
Many scholars have attempted to generalize the Erdős matching conjecture to the transversal version in hypergraph systems. Matsumoto and Tokushige [119] generalized the Erdős–Ko–Rado theorem to two families.
@@
It is natural to extend the Erdős–Ko–Rado theorem to more than two hypergraphs. For this, Aharoni and Howard [6] and Huang, Loh and Sudakov [76] proposed the following conjecture independently.
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{t}\}$ be a system of $k$-graphs on a common vertex set of size $n$. If $|E(H_{i})|>\max\left\{{\binom{n}{k}}-{\binom{n-t+1}{k}},{\binom{kt-1}{k}}\right\}$ for all $1\leq i\leq t$, then $\mathcal{H}$ has a transversal $t$-matching.
When $k=2$, this conjecture was confirmed by Meshulam (see [6]). Indeed, the case $k=2$ of this conjecture follows directly from an earlier result established by Akiyama and Frankl [8], as restated in [55]. Huang, Loh and Sudakov [76] proved that Conjecture 7.10 holds for $n>3k^{2}t$. Keller and Lifshitz [90] proved that Conjecture 7.10 holds when $n\geq f(t)k$ for some large constant $f(t)$ which only depends on $t$, and this was further improved the bound on $n$ to $12kt\log(e^{2}t)$ by Frankl and Kupavskii [58]. Lu, Wang and Yu [115] improved this bound to $n>2kt$ for sufficiently large $t$. By using the sharp threshold techniques developed in [85], Keevash, Lifshitz, Long and Minzer [84] proved a more general result with $n=\Omega(kt)$, which is the optimal order of magnitude. Kupavskii [104] gave the concrete dependencies on the parameters by showing the conjecture holds for $n>3ekt$ with $t>10^{7}$. Recently, Gao, Lu, Ma and Yu [59] showed that Conjecture 7.10 is true when $k=3$ and $n$ is sufficiently large.
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{t}\}$ be a $k$-partite $k$-graph system with $n$ vertices in each partition class. If $|E(H_{i})|\geq(t-1)n^{k-1}$ for all $i\in[t]$, then $\mathcal{H}$ admits a transversal $t$-matching.
Aharoni and Howard proved that Conjecture 7.11 holds when $t=2$ [6] or $k\leq 3$ [7]. Lu and Yu [116] proved Conjecture 7.11 for n>3(k1)(t1)n>3(k-1)(t-1). Kiselev and Kupavskii [94] showed that this conjecture is true for $t\geq t_{0}$ where $t_{0}$ is a constant.
The degree conditions for the existence of large matchings in uniform hypergraphs have been extensively studied. Kühn, Osthus and Treglown [103] (also see Khan [91]) and Khan [92] determined the minimum $1$-degree condition for an $n$-vertex 33-graph and $4$-graph to contain a perfect matching, respectively. Lu, Yu and Yuan [117] and Lu, Wang and Yu [114] proved the transversal version of the above results.
@@
Let $n$ be a sufficiently large integer and $k\in\{3,4\}$ with $k|n$. If $\mathcal{H}=\{H_{1},\ldots,H_{\frac{n}{k}}\}$ is a $k$-graph system of order $n$ with $\delta_{1}(\mathcal{H})>\binom{n-1}{k}-\binom{(k-1)n/k}{k}$, then $\mathcal{H}$ contains a transversal perfect matching.
Recently, Lu and Wang [113] determined the minimum $1$-degree condition that guarantees a $3$-partite 33-graph system admits transversal perfect matchings. This generalized a result of Lo and Markström [112] on the $1$-degree threshold for the existence of perfect matchings in $3$-partite 33-graphs. Moreover, in their proof, they used fractional transversal matching theory obtained by Aharoni, Holzman and Jiang [5].
@@
Aharoni, Holzman and Jiang [5] studied the existence of rainbow fractional matchings in a $k$-graph system.
@@
Let $k\geq 2$ be an integer, and let $n$ be a positive rational number. Assume that $\mathcal{H}=\{H_{1},\ldots,H_{\lceil kn\rceil}\}$ is a $k$-graph system with $\nu^{*}(H_{i})\geq n$ for all $i\in[\lceil kn\rceil]$. Then $\mathcal{H}$ has a rainbow fractional matching of size $n.$
When the $k$-graph is $k$-partite and $n$ is an integer, the number of graphs in $\mathcal{H}$ required in Theorem 7.13 goes down from $kn$ to knk+1.kn-k+1. Theorem 7.13 provided a fractional version of the corresponding problem concerning rainbow matchings, which was solved by Drisko [44] and by Aharoni and Berger [1] for bipartite graphs (the rainbow matching problem remains open for general graphs as well as for $k$-partite hypergraphs when $k>2$).
Alon, Frankl, Huang, Rödl, Ruciński and Sudakov [9] established the relationship between perfect matching and perfect fractional matching by showing that every $n$-vertex $k$-graph $H$ with $\delta_{d}(H)\geq(\max\{c_{k,d},1/2\}+o(1))\binom{n-d}{k-d}$ has a perfect matching, and this condition is asymptotically best possible. Cheng, Han, Wang and Wang [28] proved that the $d$-degree condition that ensures the existence of a perfect matching in a $k$-graph also guarantees the existence of a transversal perfect matching in a $k$-graph system for $d\in[k-1]$.
@@
For every $\varepsilon>0$ and integer $d\in[k-1]$, there exists $n_{0}\in\mathbb{N}$, such that the following holds for all integers $n\geq n_{0}$ and $n\in k\mathbb{N}$. Every $n$-vertex $k$-graph system $\mathcal{H}=\{H_{1},\ldots,H_{\frac{n}{k}}\}$ with $\delta_{d}(\mathcal{H})\geq(\max\{c_{k,d},\frac{1}{2}\}+\varepsilon)\binom{n-d}{k-d}$ contains a transversal perfect matching.
Lang [105] gave a common generalization of the above theorem. Let $G_{1}$ and $G_{2}$ be $k$-graphs on the same vertex set $V$, and let $F$ be a $k$-graph. A color covering homomorphism from $F$ to $(G_{1},G_{2})$ is a function $\phi:V(F)\rightarrow V$ with an edge $f\in E(F)$ such that $\phi(f)\in E(G_{1})$ and $\phi$ is a homomorphism from $F^{\prime}$ to $G_{2}$, where $F^{\prime}$ is obtained from $F$ by deleting $f$ from its edge set. We define $\delta_{d}^{C_{F}}$ to be the smallest real number $\delta$ such that for all $\alpha>0$ and $n$ large enough any two $k$-graphs $G_{1},G_{2}$ on the same $n$ vertices with $\delta_{d}(G_{1}),\delta_{d}(G_{2})\geq(\delta+\alpha)\binom{n-d}{k-d}$ admit a color covering homomorphism from $F$ to $(G_{1},G_{2})$. Denote by $\delta_{d}^{F}$ the smallest real number $\delta$ such that for all $\alpha>0$ and sufficiently large $n$, every $n$-vertex $k$-graph $G$ with $\delta_{d}(G)\geq(\delta+\alpha)\binom{n-d}{k-d}$ contains a copy of $F$-factor.
Theorem 7.15 ([105]).
Let $F$ be a $k$-graph with $r$ vertices and $t$ edges. For every $\varepsilon>0$ and integer $d\in[k-1]$, there exists $n_{0}\in\mathbb{N}$ such that the following holds for all integers $n\geq n_{0}$. Every $rn$-vertex $k$-graph system $\mathcal{H}=\{H_{1},\ldots,H_{tn}\}$ with $\delta_{d}(\mathcal{H})\geq\left(\max\left\{\delta_{d}^{C_{F}},\delta_{d}^{F}\right\}+\varepsilon\right)\binom{n-d}{k-d}$ contains a transversal $F$-factor.
This is essentially a hypergraph version of the work of Montgomery, Müyesser and Pehova (see Theorem 3.12). As an application of Theorem 7.15, Lang [105] deduced the minimum degree condition that guarantees the existence of transversal $K_{4}^{3}$-factors in a $3$-graph or transversal power of a Hamilton cycles in a $k$-graph, where $K_{4}^{3}$ is the complete $3$-graph on $4$ vertices.
@@ 8 Further results
  • The rainbow Turán problem can be viewed as a special case of Turán problems in graph systems. Let $H$ be a graph. The rainbow Turán number of $H$, denoted by $\mathrm{ex}^{*}(n,H)$, is defined to be the maximum number of edges in a properly edge-colored $n$-vertex graph $G$ with no rainbow copies of $H$. Notice that this properly edge-colored graph $G$ can be viewed as a system $\mathcal{G}=\{G_{1},\ldots,G_{\chi(G)}\}$ of graphs where each $G_{i}$ is the graph consisting of edges with color $i.$ Then finding a rainbow copy of $H$ in $G$ is equivalent to finding a transversal copy of $H$ inside $\mathcal{G}$. The rainbow Turán number of some classical families of graphs were studied by Keevash, Mubayi, Sudakov and Verstraëte [86]. See more results in [17, 38, 79, 93].
  • In section 3, we mainly survey results on the minimum degree condition for transversal spanning structures in graph systems. In fact, some studies have considered the problem of embedding transversal non-spanning structures in a graph system. Li, Li and Li [108] studied the existence of transversal cliques in graph systems under Dirac-type conditions. Falgas-Ravry, Markström and Räty [51] studied the minimum degree version of Problem 2.2, and they found that the extremal behavior in the minimum degree condition differs strikingly from that seen in the density condition, with discrete jumps as opposed to continuous transitions.
  • The adjacency matrix of $G$ is $A(G)=(a_{ij})$, where $a_{xy}=1$ if $xy\in E(G)$ and 0 otherwise. The largest eigenvalue of $A(G)$ is called the spectral radius of $G$. The maximum spectral radius of graphs in a graph system that forbids transversal perfect matchings, Hamilton paths or Hamilton cycles was studied in [64, 73, 143].
  • One may consider the problem of embedding a rainbow graph $H$ in a graph system $\mathcal{G}$ provided that $H$ and all graphs in $\mathcal{G}$ belong to the same class. For example, Aharoni, Briggs, Holzman and Jiang [2] proved that every family of $2\left\lceil n/2\right\rceil-1$ odd cycles of order $n$ contains a rainbow odd cycle. For more results, we refer the reader to [43, 63, 107].
  • Another classical and closely related concept is the independent transversal. Given a graph $G$ and a partition of its vertex set, an independent transversal is an independent set of $G$ that contains one vertex from each partition class. It also concerns selecting one element from each of several sets while satisfying a global structural constraint. This is a very general notion, many combinatorial problems can be formulated by asking whether a certain graph with a certain vertex partition has an independent transversal. For more related results, see [23, 37, 62, 70, 71, 72, 135, 142].
@@ 9 Conclusion
Although we have made an effort to provide a systematic coverage of recent results in graph systems, it is inevitable that some topics were left out of this survey due to constraints of space and time. Nevertheless, we would like to believe that we have presented enough results demonstrating how one can generalize known results in extremal graph theory to transversal versions. Therefore we hope this survey will inspire further research in this area.
10 Acknowledgement
We would like to express our gratitude to the anonymous reviewer for his/her valuable comments that greatly improved the presentation of this paper. We also thank Stijn Cambie, Yangyang Cheng, Jie Hu and Richard Lang for helpful comments and valuable suggestions.